# Do Current Language Models Support Code Intelligence for R Programming Language?

### v2

@@ 4.1.1. Code Summarization
Table 2 presents the BLEU-4 scores for code summarization, for all languages. The Combinedsix represents the results when we combined all six languages from CodeSearchNet. Row Rcombine shows the dataset that we combine both Tidy-verse and Base R datasets.
@@
As shown in Table 2, for all the six languages, JavaScript (JS), Go, Java, PHP, Ruby, and Python, CodeT5 has the best performance. For some of the languages, UniXCoder has close scores, all being higher than CodeBERT and GraphCodeBERT. When the models are tested on RCombine, we observe that UniXCoder has the highest performance, being $18.43$ followed by CodeT5 BLEU-4 score of $15.42$. The CodeBERT and GraphCodeBERT both have lower score in the range of $11$ BLEU-4 scores.
It is worth noting that GraphCodeBERT (Guo et al., 2020) is primarily designed to model the structural relationships in code (e.g., how different components of code like functions, variables, and classes are related). This is beneficial for tasks that rely heavily on understanding the syntactic and semantic structure of code, such as code completion or defect detection. For summarization, where the goal is to generate natural language descriptions from code, a model needs to excel in both code understanding and natural language generation. Though we tokenizaed the R dataset using the same tool, Tree-sitter, GraphCodeBERT continues to struggle for R. A possible reason could be that R, as a programming language, is fundamentally different from other languages. This difference means that R does not share the same expressions or patterns, making it challenging for the model to fit the data effectively during training. So this might be a culprit for low performance of GraphCodeBERT on R.
@@ 4.1.2. Method Name Prediction
We study the performance of fine-tuning all models using RTidy, RBase and Rcombine. As we did for code summarization, we re-trained the baseline models, to get performance numbers for each of the six programing language in the codeSearchNet (Husain et al., 2019) dataset and report all the results. Table 3 shows the F1-score of all the all the programming languages for method name prediction. GraphCodeBERT outperforms CodeBERT, UniXCoder and CodeT5 in all languages except Go which GraphCodeBERT and CodeBERT show same scores. The performance of GraphCodeBERT demonstrate that GraphCodeBERT’s graph-based approach is more effective than the other three base line models in handling method name prediction. On the other hand, Unixcoder does not continue to maintain its performance in code summarization, instead, UniXCoder shows its struggle for method name prediction, being the lowest in JS, PHP, RTidy, RBase and Rcombine. UniXcoder is designed to handle multiple code intelligence tasks in a unified manner, shows varied results across the languages. UniXcoder’s scores lower than the other models, is potentially due to its broader generalization approach, which may not specialize as effectively in code understanding. CodeT5, likewise, did not continue to maintain its performance in this task, with CodeT5 being the worst performer in the other four languages: Go, Java, Python and Ruby.
The R language datasets present significant challenges for all models. The scores are around $0.25$ for most models and around $0.32$ for the best model (i.e., GraphCodeBERT). Interestingly, for this task, there is no difference between the RTidy, RBase and Rcombine; and all R datasets have exact or very close scores. This indicated the need to handle R’s domain-specific syntax and statistical focus by further adaptation or specialized pre-training models that works for R.
@@ 4.2.1. Code Summarization
Table 4 represents the results for code summarization under multi-lingual fine-tuning. In this setting, the performance of the baseline models contradicted with previous findings (Ahmed and Devanbu, 2022), where multi-lingual fine-tuning typically improved model performance. For UniXcoder, and CodeT5, multilingual fine-tuning actually leads to performance degradation. UniXcoder experience a drop in performance for most programming languages, and CodeT5 shows slightly performance degradation only in Ruby, with significant drop in all other languages, indicating that the effectiveness of multilingual fine-tuning varies depending on the model. This drop in performance is also reported in previous works (Wang et al., 2023), and dropped even further with the addition of R language to the dataset. Specifically, for CodeBERT, the BLEU-4 scores increased for JavaScript, Go, and Ruby by $7\%$, $1.4\%$, and $28\%$ respectfully, the improvement is significant for JavaScript and Ruby, showing a modest improvement across these languages. However, other languages exhibited mixed results. For example, PHP and Java received almost no change, and for R language, the combined R dataset (RCombine) demonstrated a $39\%$ performance drop compared to mono-lingual setting.
@@
When examining GraphCodeBERT, there was a consistent trend of performance improvement across all languages except for combined R dataset (RCombine), which is decreased by $1.4$ BLEU-4 score. All six languages of CodeSearchNet data have about one BLEU score improvement, except Ruby which jumps to $14.98$ from $11.76$, seeing the highest increase. The performance on R datasets was more variable, with RBase seeing a notable increase of $25\%$ (p-value equal to 0.003), while RTidy only increases $7\%$ (with p-value equal to 0.766 which is not significant). Here we can see that although GraphCodeBERT performed well under multi-lingual fine-tuning, the combination of RTidy and RBase would hurt the effects of fine-tuning while testing them, as shown in the results of RCombine.
@@
In multi-lingual setting, UniXCoder (Guo et al., 2022) on the other hand, has reduced the BLEU scores for all languages, dropping the scores slightly. Particularly for RCombine, while it had previously demonstrated strong performance in monolingual settings, its the performance dropped by 22% (with p-value less than 0.001). For RBase and RTidy, the performance dropped slightly. This observation again highlights the negative effects from combining RTidy and RBase datasets during testing.
@@
Lastly, CodeT5 demonstrated very low performance on all languages, except JavaScript. This performance drop is more noticeable for R datasets, where the BLEU scores are around $2$. The low scores could be attributed to CodeT5’s text-to-code architecture, which may not generalize as well in multi-lingual fine-tuning. While CodeT5 performed well in the monolingual setting, its performance dipped significantly during fine-tuning on the R datasets. This points to a potential limitation in its ability to handle diverse programming paradigms across different languages. It is worth pointing out that CodeT5 also took the most time to fine-tune, and during training we observed that the model had a hard time converging to the point where the model showed a drop in performance on the dev set. Combined with the performance of other base line models, this is a good indication that pre-trained models such as CodeT5 have some shortcomings in learning a language such as R, which has its own unique syntax and structure, specially for RCombine.
@@
To evaluate the significance of the differences in model performance, we employed the Wilcoxon signed-rank test, a non-parametric statistical test commonly used for comparing paired samples. Wilcoxon signed-rank test was chosen due to its suitability for comparing two related samples, making it ideal for performance comparisons across different models on the same dataset. The p-value is computed by applying the Wilcoxon signed test on mono-lingual fine-tuning result and multi-lingual fine-tuning for code summarization and method name prediction. Under this setting the p-value represents whether the observed differences in performance are statistically significant or not. To control for the risk of false positives due to multiple comparisons, we applied the Benjamini-Hochberg (B-H) correction following previous study (Ahmed and Devanbu, 2022). The B-H correction is a widely used method for controlling the false discovery rate (FDR), which refers to the expected proportion of incorrectly rejected null hypotheses (false positives) among all rejected hypotheses. By adjusting p-values obtained from the Wilcoxon signed-rank test, the B-H correction allows us to maintain a balance between discovering true effects and limiting the inclusion of false positives when performing multiple comparisons.
@@
The p-values in Table 4 that are $<0.05$ and show the numbers are statistically significant from mono-lingual setting, are shown with an asterisk (*). In the table, some results are not significant based on their p-value. These include JavaScript for UniXCoder, Go for CodeBERT, PHP for CodeBERT, RTidy for GraphCodeBERT and UniXCoder, and RBase for UniXCoder, these higher p-values indicate that the differences in BLEU scores for these models are not statistically significant. Most results though, show a significant difference for R.
@@ 4.2.2. Method Name Prediction
Table 5. Result of method name prediction for RQ2. Best results in each row are bold and best results in each column are underlined. The p-values compared to mono-lingual setting are shown for each model. P-values smaller than $0.05$ show a statistically significant difference and are marked with an asterisk (*).
Table 5 shows the result of all the baselines for the six programming languages, R styles, and combined R dataset on method name prediction under multilingual training. We observe significant drop in performance for all four models and the six languges of CodeSearchNet dataset (the p-value is less than 0.001 for nearly all of models and language). Multilingual training generally provides models with a broader understanding of diverse languages by exposing them to a combined dataset, which helps in cross-linguistic generalization but may dilute performance on individual languages due to the need to balance between multiple linguistic patterns. For all six languages of the CodeSearchNet dataset, and for all models, the performance has dropped to $0.10s$ with only a few close to $0.2$.
In examining the R styles, especially RTidy and RBase, the results under multilingual training are notably different. GraphCodeBERT performs surprisingly well in these cases, boosting the performance; particularly under multilingual training, where it scores $0.38$ for RTidy and $0.41$ for RBase, compared to $0.31$ and $0.32$ in mono-lingual setting, which is statistically significant. This means the performance is improved by $22.58\%$ and $31.25\%$ for each of the RTidy and RBase, respectively. However, the performance of CodeBERT and UniXCoder drop slightly under multi-lingual fine-tuning, but the p-value still shows that the change is significant since it is less than 0.05. For RCombine, the performance is dropped slightly in multi-lingual setting for all models, suggesting that combining the two R styles for method name prediction is not helpful, though separate styles would benefit in multi-lingual setting.
For RCombine, RTidy, and RBase, all p-values are less than $0.05$ (marked with an asterisk), indicating statistically significant differences compared to the monolingual setting, with one exception. The exception is UniXCoder’s performance on RTidy, where the p-value of $0.227$ suggests no statistically significant difference. Notably, GraphCodeBERT consistently shows the best performance across all R datasets, with F1-scores of $0.28$, $0.38$, and $0.41$ for RCombine, RTidy, and RBase respectively, outperforming all other languages. These results are all statistically significant ($p<0.05$).
@@
CodeT5, built on the T5 architecture, frames code-related tasks as text-to-text problems. This approach allows for a unified model capable of handling various software engineering tasks, including both code summarization and method name prediction. Our results show that each model’s architecture influences its performance differently across tasks. For both tasks, we found that CodeT5’s specific architecture is experiencing significant drop in performance with drop over $50$% and p-value less than $0.001$. This suggests that a unified framework that converts language problem into a text-to-text format will extraordinarily affect the performance of the model under multi-lingual training. It is worth pointing out that CodeT5 also has an early stop mechanism, which will automatically stop fine-tuning if the training results do not improve in relation to a specific round to reduce the fine-tuning time. When fine-tuning under multi-lingual setup, the model takes much longer than mono-lingual fine-tuning, and more epoch to stop. This also suggests that converting the problem to a text-text problem under multi-lingual fine-tuning poses extra challenge for the model to convergence.
@@ 4.3.1. Code Summarization
Table 6 reports the results of code summarization using cross-project training across CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5. In cross-project training, data elements from one project are strictly allocated to either training, testing, or validation sets, ensuring that the model is tested on entirely unseen data from different projects.
@@
These results demonstrate that UniXCoder mostly delivers superior performance across different R programming styles in a cross-project training scenario, particularly excelling in Base R tasks. CodeT5 also shows competitive results, particularly in the Tidy-verse style. Though UniXCoder and CodeT5 have highest scores in the cross-project setting, their performance has dropped significantly, more than $11$ BLEU-4 scores compared to Table 2. Similar to RQ1, Base R and the combined version have higher scores compared to Tidy-verse. This finding align with the previous findings of intra- vs cross project setting (Zhang et al., 2016a), where intra-project outperforms cross-project setting in most cases.
In contrast, the intra-project results, where the models are tested on the same project they were trained on, demonstrate higher performance. CodeBERT, for example, achieves a significant improvement on the RBase dataset, scoring $31.18$, an improvement from $19.55$ BLEU-4. We observe the highest improvement for CodeBERT followed by CodeT5 compared to other models. Among the two R styles and RCombined, RBase scores increase more than the others.
Overall, the intra-project results underscore the importance of project-specific training, as models tend to perform significantly better when tested on data with a similar distribution. This is particularly evident from the large boosts in the models’ performances from cross-project to intra-project scenarios. GraphCodeBERT, however, while still improving, does not gain as much, suggesting that it may struggle more with the nuances of project-specific contexts. This is aligned with our previous findings in RQ1, where GraphCodeBERT had the lowest score among all models. CodeT5 and UniXCoder, while performing well across both settings, show a slightly smaller margin of improvement, indicating a more consistent but less specialized approach to the task.

### v3

@@ 4.1.1. Code Summarization
Table 2 presents the BLEU-4 scores for code summarization, for all languages. The Combinedsix represents the results when we combined all of six languages from CodeSearchNet. Row Rcombine shows the dataset that we combine both Tidy-verse and Base R datasets.
@@
As shown in Table 2, for most of the six languages, JavaScript (JS), Go, Java, PHP, Ruby, and Python, CodeT5 has the best performance. For some of the languages, UniXCoder has close scores, all being higher than CodeBERT and GraphCodeBERT. When the models are tested on RCombine, we observe that UniXCoder has the highest performance, being $18.43$, followed by CodeT5 BLEU-4 score of $15.42$. CodeBERT and GraphCodeBERT both have lower scores in the range of $11$ to $12$ BLEU-4 scores.
It is worth noting that GraphCodeBERT (Guo et al., 2020) is primarily designed to model the structural relationships in code (e.g., how different components of code, like functions, variables, and classes, are related). This is beneficial for tasks that rely heavily on understanding the syntactic and semantic structure of code, such as code completion or defect detection. For summarization, where the goal is to generate natural language descriptions from code, a model needs to excel in both code understanding and natural language generation. Though we tokenized the R dataset using the same tool, Tree-sitter, GraphCodeBERT continues to struggle with R. One possible reason could be that R, as a programming language, is fundamentally different from other languages. This difference means that R does not share the same expressions or patterns, making it challenging for the model to learn the data effectively during training, which might be a culprit for the low performance of GraphCodeBERT on R.
@@ 4.1.2. Method Name Prediction
We study the performance of fine-tuning all models using RTidy, RBase, and Rcombine. As we did for code summarization, we re-trained the baseline models to get performance numbers for each of the six programming languages in the codeSearchNet (Husain et al., 2019) dataset and report all the results. Table 3 shows the F1 scores for all programming languages for method name prediction. GraphCodeBERT outperforms CodeBERT, UniXCoder, and CodeT5 in all languages except Go, for which GraphCodeBERT and CodeBERT achieve the same score. The performance of GraphCodeBERT demonstrates that GraphCodeBERT’s graph-based approach is more effective than the other three baseline models in handling method name prediction. On the other hand, UniXCoder does not continue to maintain its performance in code summarization, instead, UniXCoder shows its struggle for method name prediction, being the lowest in JS, PHP, RTidy, and RBase. UniXCoder, which is designed to handle multiple code-intelligence tasks in a unified manner, shows varied results across languages. Its lower scores compared with those of the other models may stem from its broader generalization approach, which may not specialize as effectively in code understanding. Likewise, CodeT5 failed to maintain strong performance on this task and was the worst performer in the four remaining languages—Go, Java, Python, and Ruby.
The R language datasets present significant challenges for all models. The scores are around $0.25$ for most models and around $0.32$ for the best model (i.e., GraphCodeBERT). Interestingly, for this task, there is no difference between the RTidy, RBase, and Rcombine; and all R datasets have exact or very close scores. This indicated the need to handle R’s domain-specific syntax and statistical focus by further adaptation or specialized pre-training models that work for R.
@@ 4.2.1. Code Summarization
Table 4 represents the results for code summarization under multi-lingual fine-tuning. In this setting, the performance of the baseline models contradicted with previous findings (Ahmed and Devanbu, 2022), where multi-lingual fine-tuning typically improved model performance. For UniXCoder, and CodeT5, multilingual fine-tuning actually leads to performance degradation. UniXCoder experiences a drop in performance for most programming languages, and CodeT5 shows slight performance degradation only in Ruby, with a significant drop in all other languages, indicating that the effectiveness of multilingual fine-tuning varies depending on the model. This drop in performance is also reported in previous works (Wang et al., 2023), and dropped even further with the addition of R language to the dataset. Specifically, for CodeBERT, the BLEU-4 scores increased for JavaScript, Go, and Ruby by $7\%$, $1.4\%$, and $28\%$, respectively. The improvement is significant for JavaScript and Ruby, showing a modest improvement across these languages. However, other languages exhibited mixed results. For example, PHP and Java showed almost no change, and for the R language, the combined R dataset (RCombine) demonstrated a $39\%$ performance drop compared to the monolingual setting.
@@
When examining GraphCodeBERT, there was a consistent trend of performance improvement across all languages except for combined R dataset (RCombine), which is decreased by $1.4$ BLEU-4 score. All six languages of CodeSearchNet data have about one BLEU score improvement, except Ruby, which jumps to $14.98$ from $11.76$, seeing the highest increase. The performance on R datasets was more variable, with RBase seeing a notable increase of $25\%$ (p-value equal to 0.003), while RTidy only increases $7\%$ (with p-value equal to 0.766, which is not significant). Here we can see that although GraphCodeBERT performed well under multi-lingual fine-tuning, the combination of RTidy and RBase would hurt the effects of fine-tuning while testing them, as shown in the results of RCombine.
@@
In multi-lingual setting, UniXCoder (Guo et al., 2022) on the other hand, shows reduced BLEU scores for all languages, albeit slightly. Particularly for RCombine, while it had previously demonstrated strong performance in monolingual settings, its performance dropped by 22% (with p-value less than 0.001). For RBase and RTidy, the performance dropped slightly. This observation again highlights the negative effects of combining RTidy and RBase datasets during testing.
@@
Lastly, CodeT5 demonstrated very low performance on all languages, except JavaScript. This performance drop is more noticeable for R datasets, where the BLEU scores are around $2$. The low scores could be attributed to CodeT5’s text-to-code architecture, which may not generalize as well in multi-lingual fine-tuning. While CodeT5 performed well in the monolingual setting, its performance dipped significantly during fine-tuning on the R datasets. This points to a potential limitation in its ability to handle diverse programming paradigms across different languages. It is worth pointing out that CodeT5 also took the most time to fine-tune, and during training, we observed that the model had a hard time converging to the point where the model showed a drop in performance on the dev set. Combined with the performance of other baseline models, this is a good indication that pre-trained models such as CodeT5 have some shortcomings in learning a language such as R, which has its own unique syntax and structure, especially for RCombine.
@@
To evaluate the significance of the differences in model performance, we employed the Wilcoxon signed-rank test, a non-parametric statistical test commonly used for comparing paired samples. Wilcoxon signed-rank test was chosen due to its suitability for comparing two related samples, making it ideal for performance comparisons across different models on the same dataset. The p-value is computed by applying the Wilcoxon signed test on mono-lingual fine-tuning result and multi-lingual fine-tuning for code summarization and method name prediction. Under this setting the p-value represents whether the observed differences in performance are statistically significant or not. To control for the risk of false positives due to multiple comparisons, we applied the Benjamini-Hochberg (B-H) correction following a previous study (Ahmed and Devanbu, 2022). The B-H correction is a widely used method for controlling the false discovery rate (FDR), which refers to the expected proportion of incorrectly rejected null hypotheses (false positives) among all rejected hypotheses. By adjusting p-values obtained from the Wilcoxon signed-rank test, the B-H correction allows us to maintain a balance between discovering true effects and limiting the inclusion of false positives when performing multiple comparisons.
@@
The p-values in Table 4 that are $<0.05$ and show the numbers are statistically significant from mono-lingual setting, are shown with an asterisk (*). In the table, some results are not significant based on their p-value. These include JavaScript for UniXCoder, Go for CodeBERT, PHP for CodeBERT, RTidy for GraphCodeBERT and UniXCoder, and RBase for UniXCoder, these higher p-values indicate that the differences in BLEU scores for these models are not statistically significant. Most results, though, show a significant difference for R.
@@ 4.2.2. Method Name Prediction
Table 5. Result of method name prediction for RQ2. Best results in each row are bold, and best results in each column are underlined. The p-values compared to mono-lingual setting are shown for each model. P-values smaller than $0.05$ show a statistically significant difference and are marked with an asterisk (*).
Table 5 shows the result of all the baselines for the six programming languages, R styles, and combined R dataset on method name prediction under multilingual training. We observe a significant drop in performance for all four models and the six languages of the CodeSearchNet dataset (the p-value is less than 0.001 for nearly all of the models and languages). Multilingual training generally provides models with a broader understanding of diverse languages by exposing them to a combined dataset, which helps in cross-linguistic generalization but may dilute performance on individual languages due to the need to balance between multiple linguistic patterns. For all six languages of the CodeSearchNet dataset, and for all models, the performance has dropped to $0.10s$ with only a few close to $0.2$.
In examining the R styles, especially RTidy and RBase, the results under multilingual training are notably different. GraphCodeBERT performs surprisingly well in these cases, boosting the performance; particularly under multilingual training, where it scores $0.38$ for RTidy and $0.41$ for RBase, compared to $0.31$ and $0.32$ in a mono-lingual setting, which is statistically significant. This means the performance is improved by $22.58\%$ and $31.25\%$ for each of the RTidy and RBase, respectively. However, the performance of CodeBERT and UniXCoder drops slightly under multi-lingual fine-tuning, but the p-value still shows that the change is significant since it is less than 0.05. For RCombine, the performance is dropped slightly in a multi-lingual setting for all models, suggesting that combining the two R styles for method name prediction is not helpful, though separate styles would benefit in a multi-lingual setting.
For RCombine, RTidy, and RBase, all p-values are less than $0.05$ (marked with an asterisk), indicating statistically significant differences compared to the monolingual setting, with one exception. The exception is UniXCoder’s performance on RTidy, where the p-value of $0.227$ suggests no statistically significant difference. Notably, GraphCodeBERT consistently shows the best performance across all R datasets, with F1-scores of $0.28$, $0.38$, and $0.41$ for RCombine, RTidy, and RBase, respectively, outperforming all other languages. These results are all statistically significant ($p<0.05$).
@@
CodeT5, built on the T5 architecture, frames code-related tasks as text-to-text problems. This approach allows for a unified model capable of handling various software engineering tasks, including both code summarization and method name prediction. Our results show that each model’s architecture influences its performance differently across tasks. For both tasks, we found that CodeT5’s specific architecture is experiencing a significant drop in performance with a drop of over $50$% and a p-value less than $0.001$. This suggests that a unified framework that converts language problems into a text-to-text format will extraordinarily affect the performance of the model under multi-lingual training. It is worth pointing out that CodeT5 also has an early stop mechanism, which will automatically stop fine-tuning if the training results do not improve in relation to a specific round to reduce the fine-tuning time. When fine-tuning under a multi-lingual setup, the model takes much longer than mono-lingual fine-tuning, and more epochs to stop. This also suggests that converting the problem to a text-text problem under multi-lingual fine-tuning poses an extra challenge for the model to converge.
Although multilingual fine-tuning can incorporate knowledge from a wide array of languages, thereby exposing the model to diverse syntactic structures and potentially improving generalization, R has domain-specific characteristics—particularly statistical functions and specialized libraries—that may not align with the broader patterns learned from other programming languages. This mismatch can dilute the relevance of learned representations, preventing the model from fully capturing R’s unique semantics. In other words, while multilingual data might grant a broader contextual understanding, it risks overshadowing the specialized needs of R code, which can ultimately hinder performance when domain-specific nuances are critical, which is supported by our result in this section.
@@ 4.3.1. Code Summarization
Table 6 reports the code-summarization results under cross-project training for CodeBERT, GraphCodeBERT, UniXCoder, and CodeT5. In cross-project training, all code snippets from a given project are assigned exclusively to the training, validation, or test set, ensuring that the model is tested on data from projects it has never seen.
@@
These results demonstrate that UniXCoder mostly delivers superior performance across different R programming styles in a cross-project training scenario, particularly excelling in Base R tasks. CodeT5 also shows competitive results, particularly in the Tidy-verse style. Though UniXCoder and CodeT5 have the highest scores in the cross-project setting, their performance has dropped significantly, more than $11$ BLEU-4 scores compared to Table 2. Similar to RQ1, Base R and the combined version have higher scores compared to Tidy-verse. This finding aligns with previous work on intra- vs. cross-project settings (Zhang et al., 2016a), where intra-project outperforms cross-project settings in most cases.
In contrast, the intra-project results, where the models are tested on the same project they were trained on, demonstrate higher performance. CodeBERT, for example, achieves a significant improvement on the RBase dataset, scoring $31.18$, an improvement from $4.13$ BLEU-4. We observe the highest improvement for CodeBERT, followed by CodeT5, compared to other models. Among the two R styles and RCombined, RBase scores increase more than the others.
Overall, the intra-project results underscore the importance of project-specific training, as models tend to perform significantly better when tested on data with a similar distribution. This is particularly evident from the large boosts in the models’ performances from cross-project to intra-project scenarios. GraphCodeBERT, however, although it still improves, does not gain as much, suggesting that it may struggle more with the nuances of project-specific contexts. This aligns with our previous findings from RQ1, where GraphCodeBERT had the lowest score among all models. CodeT5 and UniXCoder, while performing well across both settings, show a slightly smaller margin of improvement, indicating a more consistent but less specialized approach to the task.